- you want to buy a brand-new car. Its price is $49,000 and the dealer’s special financing offers you 2.75% APR. What will be your monthly payment if you want to pay the car loan off in 6 years?

How much interest money will you save if you make more payments and pay the loan off in 3 years?
Answer #1
1) Monthly Payment = Loan amount / Present value of annuity of 1
= $ 49,000.00 / 66.3038053
= $       739.02
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.002292)^-72)/0.002292 i = 2.75%/12 = 0.002292
= 66.3038053 n = 6*12 = 72
2) Interest saved $   2,104.48
Working:
# 1 Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.002292)^-36)/0.002292 i = 2.75%/12 = 0.002292
= 34.5170816 n = 3*12 = 36
# 2 Monthly Payment = Loan amount / Present value of annuity of 1
= $ 49,000.00 / 34.5170816
= $   1,419.59
# 3 6 Years 3 Years Difference
Monthly Payment $       739.02 $   1,419.59
Months 72 36
Total repayments $ 53,209.62 $ 51,105.13
Loan $ 49,000.00 $ 49,000.00
Interest $   4,209.62 $   2,105.13 $ 2,104.48